OpenStudy (emilyjones284):

Find the horizontal and vertical components of the vector with given length and direction, and write the vector in terms of the vectors i and j. absolute value of v =50, theta=120 degrees

ganeshie8 (ganeshie8):

x component : \(r \cos\theta\) y component : \(r \sin\theta\)

ganeshie8 (ganeshie8):

\(r\) is the magnitude of vector

OpenStudy (doc.brown):



OpenStudy (emilyjones284):

wait so how do you get the answer?

OpenStudy (emilyjones284):

is that the answer?

OpenStudy (doc.brown):

I don't know why my \(\theta\)'s came out like Q's in the drawing.

OpenStudy (emilyjones284):

so is that the answer or...

OpenStudy (emilyjones284):

pleease help me

OpenStudy (whpalmer4):

Here's how you get the x and y components of any vector: x component is \(r \cos \theta\) y component is \(r \sin \theta\) In your case, they've referred to \(r\) as \(v\). Don't be fooled :-)

OpenStudy (whpalmer4):

Note that if the angle is 0, the vector points to the right along the x-axis, and the x component is the entire vector length, because you are multiplying by \(\cos 0 = 1\). The y component is 0, because you are multiplying by \(\sin 0 = 0\). If \(\theta\) is \(90^\circ\) or \(\pi/2\) radians, the vector is entirely y component, with no x component. Usually, you'll end up with an angle other than a multiple of \(90^\circ\), and there will be both x and y components.

OpenStudy (emilyjones284):

thank you!

OpenStudy (emilyjones284):

That is very helpful (:

OpenStudy (whpalmer4):

You're welcome! It's really pretty simple, perhaps that's what confused you :-)

OpenStudy (emilyjones284):

Yeah I think too much about some things haha

OpenStudy (whpalmer4):

It might help if you think about the unit circle and the values of cos and sin as theta varies. When your vector points to the right, down the x axis, all of the magnitude goes into the x component. As the angle opens, less of it goes to the right and more goes up in the y component, until at the top of the unit circle, there is no x component, it's all y component. Also, if you've done parametric equations, the parametric equations for a circle should remind you of the decomposition of a vector, and vice versa. http://www.mathopenref.com/coordparamcircle.html

OpenStudy (emilyjones284):

okay thank you (: